last month a store manager decided to decrease the prices of all items by 10%. this month he increased the prices by 10%. What would be the price for a pair of pants that had cost $750 before prices were decreased last month??

Please include your solution... thank you in advance!! :)

Found 4 solutions by jim_thompson5910, lynnlo, ikleyn, josgarithmetic:
Answer by jim_thompson5910(35256)

I'm assuming that the pair of pants start out at a price of $7.50

First decrease by 10%

7.50 - 10% of 7.50 = 7.50 - 0.10*7.50

7.50 - 10% of 7.50 = 7.50 - 0.75

7.50 - 10% of 7.50 = 6.75

-------------------------------------------------------
Now increase by 10%

6.75 + 10% of 6.75 = 6.75 + 0.10*6.75

6.75 + 10% of 6.75 = 6.75 + 0.675

6.75 + 10% of 6.75 = 7.425

6.75 + 10% of 6.75 = 7.43

-------------------------------------------------------
To sum things up, you

Start with 7.50

Decrease 7.50 by 10% to get 6.75

Increase 6.75 by 10% to get 7.43

So the final price is $7.43

Answer by lynnlo(4176)
Answer by ikleyn(53742)

.
last month a store manager decided to decrease the prices of all items by 10%. this month he increased the prices by 10%.
What would be the price for a pair of pants that had cost $750 before prices were decreased last month??
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


Last month the price decreased from 750 dollars to 750 - 0.1*750 = 750*(1-0.1) = 750*0.9 = 675 dollars.


This month the price increased from 675 dollars to 675 + 0.1*675 = 675*(1+0.1) = 675*1.1 = 742.50 dollars.


    
                           Notice here 
    that the price did not return to 750 dollars, the original level !



You can compose these long calculations into one short composition for two steps

      750 dollars  ----->  750*0.9*1.1 = 750*0.99 = 742.50 dollars


which gives you the same answer.


Notice this factor 0.99, which prevents the price to return to the original level ! !

Solved, with explanations.


/\/\/\/\/\/\/\/\/\/\/\/


Ignore the post by @lynnlo, since it gives wrong answer and is irrelevant.



Answer by josgarithmetic(39790)